OpenStudy (dida):

use the concept of slope to find t such that three points are collinear (-3,3) (t,-1) (8,6) i really dont understand this problem...step by step anyone?

OpenStudy (anonymous):

you need so make sure the slopes are the same

OpenStudy (anonymous):

the slope of the line between (-3,3) and (8,6) is \[\frac{6-3}{8-(-3)}\] \[=\frac{3}{11}\]

OpenStudy (anonymous):

and so you have to make sure that \[\frac{-1-3}{t-(-3)}=\frac{-4}{t+3}=\frac{3}{11}\]

OpenStudy (anonymous):

you good from there?

OpenStudy (dida):

yea and thats where i got stuck

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

cross multiply to get \[-44=3(t+3)\] how about now?

OpenStudy (dida):

yes im following

OpenStudy (anonymous):

\[-44=3t+9\] \[-53=3t\] \[t=-\frac{53}{3}\]hmm looks bad let me check

OpenStudy (dida):

yea i think thats right. i understand it way better

OpenStudy (dida):

thans!!!

OpenStudy (anonymous):

you could also find the equation for the line, then replace y by -1

OpenStudy (anonymous):

equation is \[-3 x+11 y-42 = 0\]

OpenStudy (anonymous):

let's see if it works

OpenStudy (dida):

ok

OpenStudy (anonymous):

\[-3x+11\times -1-42=0\] \[-3x=53\] yes it works

OpenStudy (anonymous):

whew i got nervous

OpenStudy (dida):

so i could use both ways right?

OpenStudy (anonymous):

sure your choice. find the line, plug in the point or make sure the slopes match up. i think first way might be shorter because in both you still have to find the slope but in second way you also have to find the equation of the line

OpenStudy (anonymous):

i mean "in the first way you still have to find the equation of the line"

OpenStudy (dida):

oh ok thanks for your help!!